OpenStudy (anonymous):

Solve log(-5x) = log(10 - 3x).

OpenStudy (anonymous):

x = 5 x = -5 x = 6 x = -8

OpenStudy (anonymous):

\[ \log(-5x) = \log(10 - 3x)\] \[e^{ (\log(-5x) )}= e^{(\log(10 - 3x))}\] \[−5x=10−3x\]

OpenStudy (anonymous):

When I took e^ to both sides the logs went away...can you solve for x now?

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

Did you even try?

OpenStudy (anonymous):

another way to view this is that log is a one to one function, so if \[\log(\text{something})=\log(\text{something else})\] then \[\text{something}=\text{something else}\]

OpenStudy (anonymous):

uhm yess i did. i added+5 to both sides

OpenStudy (anonymous):

\[−5x=10−3x \]\[−5x+5x=10−3x +5x \]\[0=10+2x \]

OpenStudy (anonymous):

Subtract 10 to both sides and divide by 2....what do you get now?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Do you know how to add subtract divide and multiply?

OpenStudy (anonymous):

dont try me

OpenStudy (anonymous):

of course i do

OpenStudy (anonymous):

You should be able to get the answer from where I left of. you only have to do two steps. subtract and divide.

OpenStudy (anonymous):

i get 0!!

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yes good job....sorry for getting at you....